### v1

@@ 6 Conclusions
The highest energy window ever probed in gamma-ray astronomy has recently been opened by LHAASO observations, reporting extended and often complex emitting regions located in the Plane of the Galaxy. The limited angular resolution of EAS arrays has so far impeded the clear identification of the nature of these objects, which represents one of the next major goals of the VHE community. Improved morphological measurements by next-generation IACTs are expected to reduce source confusion from nearby objects. At the same time, lowering the angular resolution will allow us to perform significantly better correlation studies between gamma-ray and multi-wavelength data, which are very relevant to unveil the origin of the gamma radiation. Moreover, accurate spectral analyses might provide further insights into the physical mechanisms at the origin of the radiation. Joint analysis of IACTs and LHAASO data will help in reducing spectral uncertainties, thus improving the modeling.
@@
In order to get a realistic evaluation of IACT detection prospects and required exposure times with regards to known extended gamma-ray sources, we have performed a comparative analysis of ASTRI, CTA, and LHAASO. Starting from their expected performances, we defined a method to obtain differential sensitivities towards the observation of extended sources. We found minimum detectable fluxes at the level of erg cm s for 50 hr exposure by CTA-N (at 3 TeV), erg cm s for 50 hr exposure by CTA-S (at 8 TeV), erg cm s for 50 hr exposure by ASTRI (at 3.5 TeV), and erg cm s for 1 yr exposure by LHAASO (at 100 TeV), all evaluated with disk-like sources of angular radius equal to and at zenith angle observations. We have further explored the expected IACT detection improvements for high zenith angle observations at the highest energies.
@@
Since our analysis relies on publicly available IRFs provided by each collaboration, our method to assess the sensitivity for extended sources does not include any optimization of the tools for reconstruction of the properties of primary particles, neither of dedicated background rejection methods. A precise estimation of the instrument capabilities for the detection of extended sources would require a complete 3D analysis: particularly for IACTs, the study of sub-structures on arcminute scales would permit to resolve the morphological details of many extended sources beyond the disk-like structure considered here. Moreover, we considered a single shape for the source, a uniform disk, but our methods can in principle be extended to any source morphology, including the case of asymmetric shapes for which we expect some differences as a consequence of non-uniform responses of the instruments across the FoV. Future studies will deal with the contamination of nearby sources as well as the determination of the contribution of the Galactic diffuse emission as an additional background source.
@@
Our investigation enabled us to identify the most promising extended targets, considering both the first LHAASO catalog and the HGPS, to be observed with the next generation of IACTs. We remark that the detection assessment here derived should be regarded as quite conservative, because it resulted from considering region of interests of the same extent as that announced by LHAASO; on the other hand, if the the actual source size were smaller than claimed, a detection from IACTs may require a shorter exposure than what estimated here. In particular, IACTs are expected to be sensitive to emission peaks, that might be spread out in EAS observations.
@@ Acknowledgements.
- Aharonian et al. (2021b) Aharonian, F., An, Q., Axikegu, Bai, L. X., et al. 2021b, Phys. Rev. Lett., 126, 241103
@@
- Cao et al. (2023) Cao, Z., Aharonian, F., An, Q., et al. 2023, arXiv e-prints, arXiv:2305.17030
@@
- Dong et al. (2024) Dong, X., Li, C., & He, H. 2024, arXiv e-prints, arXiv:2401.01038
@@
- H. E. S. S. Collaboration et al. (2023) H. E. S. S. Collaboration, Aharonian, F., Ait Benkhali, F., et al. 2023, A&A, 673, A148
@@
- Morlino & Celli (2021) Morlino, G. & Celli, S. 2021, MNRAS, 508, 6142
@@
Appendix A IACT worsening of IRFs with offset
@@
In the case of IACTs, extended source sensitivity is further degraded with respect to that presented in Sec. 4 if observations are carried out off-axis. Reduced performance are expected both in angular resolution and effective area. A preliminary estimation of the differential sensitivity degradation for extended source observations was performed in Ambrogi et al. (2018) for the previous configuration of CTA-South: a limited worsening was obtained for sources with radial extension below 2 deg (within a factor 2 for energies above 50 GeV). Nowadays, public IRFs are available including off-axis responses, such that more detailed evaluations can be performed. In Figs. 10, 11, and 12 we show respectively ASTRI, CTA-North, and CTA-South performances as a function of the offset angle of the observation with respect to the pointing direction of the telescope (fixed to zenith of 20 deg). For ASTRI, the impact of off-axis observations will mainly consist into a reduced effective area and enlarged angular resolution. For CTA, in addition, an increase in the energy threshold is expected.
@@
Appendix B IACT sensitivity for high zenith angle observations
IACT observations at high-zenith angles are characterized by a larger effective area at high energies at the expenses of an increased energy threshold. In fact, high-energy quasi-horizontal showers can profit of a better visibility, thanks to the larger portion of the atmosphere where they develop. On the other hand, low-energy high-zenith showers are less efficiently detected as their first interaction vertex occurs further away from the telescopes. As a result, the differential sensitivity of IACTs are affected by observations of sources at different zenith angles during the acquisition. We quantify the impact of this effect by computing the expected sensitivity for point-like observations of the CTA-N and CTA-S observatories, as shown in Fig. 14. As visible, an improvement of a factor 2 is expected above a few tens of TeV for CTA. We expect similar considerations to hold for ASTRI as well, however we cannot proceed to a proper evaluation of it since IRFs at high-zenith angle observations are not yet publicly available.
Appendix C Point-like differential sensitivities: comparison with official results and investigation of CTA subarrays

### v2

@@ 7 Conclusions
The highest energy window ever probed in gamma-ray astronomy has recently been opened by LHAASO observations, reporting extended and often complex emitting regions located in the plane of the Galaxy. The limited angular resolution of EAS arrays has so far impeded the clear identification of the nature of these objects, which represents one of the next major goals of the VHE community. Improved morphological measurements by next-generation IACTs are expected to reduce source confusion from nearby objects. At the same time, lowering the angular resolution will allow us to perform significantly better correlation studies between gamma-ray and multi-wavelength data, which are very relevant to unveil the origin of the gamma radiation. Moreover, accurate spectral analyses might provide further insights into the physical mechanisms at the origin of the radiation. Follow-up analyses with IACTs on LHAASO sources will help to improve the spectral modelling and therefore the source understanding overall.
@@
In order to get a realistic evaluation of IACT detection prospects and required exposure times with regards to known extended gamma-ray sources, we have performed a comparative analysis of ASTRI, CTA, and LHAASO. Starting from their expected performances, we defined a method to obtain differential sensitivities towards the observation of extended sources. We found minimum detectable fluxes at the level of erg cm-2 s-1 for a 50 hr exposure by CTA-N (at 3 TeV), erg cm-2 s-1 for a 50 hr exposure by CTA-S (at 8 TeV), erg cm-2 s-1 for a 50 hr exposure by ASTRI (at 3.5 TeV), and erg cm-2 s-1 for a 1 yr exposure by LHAASO (at 100 TeV); all were evaluated with disk-like sources with an angular radius equal to and at zenith angle observations. We further explored the expected IACT detection improvements for high zenith angle observations at the highest energies.
Since our analysis relies on publicly available IRFs provided by each collaboration, our method to assess the sensitivity for extended sources includes neither any optimisation of the tools for the reconstruction of the properties of primary particles, nor dedicated background rejection methods. In addition, a precise estimation of the instrument capabilities for the detection of extended sources would require a complete spectro-morphological analysis that goes beyond the simplified photon-counting approached adopted here.
@@
Moreover, we considered a single shape for the source, a uniform disk, but our methods can in principle be extended to any source morphology, including the case of asymmetric shapes for which we expect some differences as a consequence of non-uniform responses of the instruments across the FoV. For IACTs, the study of sub-structures on arc-minute scales would permit the morphological details of many extended sources to be resolved beyond the disk-like structure considered here. Future studies will deal with the contamination of nearby sources as well as the determination of the contribution of the Galactic diffuse emission as an additional background source.
@@
Our investigation enabled us to identify the most promising extended targets, considering both the first LHAASO catalogue and HGPS, that are to be observed with the next generation of IACTs. We remark that the detection assessment derived here should be regarded as rather conservative because it resulted from considering region of interests of the same extent as that announced by LHAASO; on the other hand, if the actual source size were smaller than claimed, a detection from IACTs may require a shorter exposure than what has been estimated here. In particular, IACTs are expected to be sensitive to emission peaks, which might be spread out in EAS observations.
@@ Acknowledgements.
- Aharonian et al. (2021b) Aharonian, F., An, Q., Axikegu, et al. 2021b, Chinese Physics C, 45, 025002
- Aharonian et al. (2021c) Aharonian, F., An, Q., Axikegu, Bai, L. X., et al. 2021c, Phys. Rev. Lett., 126, 241103
@@
- Albert et al. (2019) Albert, A., Alfaro, R., Ashkar, H., et al. 2019, arXiv e-prints, arXiv:1902.08429
@@
- Cao et al. (2024a) Cao, Z., Aharonian, F., An, Q., et al. 2024a, ApJS, 271, 25
- Cao et al. (2024b) Cao, Z., Aharonian, F., An, Q., et al. 2024b, Chinese Physics C, 48, 065001
@@
- Celli et al. (2024) Celli, S., Specovius, A., Menchiari, S., Mitchell, A., & Morlino, G. 2024, A&A, 686, A118
@@
- Devin et al. (2021) Devin, J., Renaud, M., Lemoine-Goumard, M., & Vasileiadis, G. 2021, A&A, 647, A68
@@
- H. E. S. S. Collaboration et al. (2011) H. E. S. S. Collaboration, Abramowski, A., Acero, F., et al. 2011, A&A, 528, A143
- H. E. S. S. Collaboration et al. (2023a) H. E. S. S. Collaboration, Aharonian, F., Ait Benkhali, F., et al. 2023a, A&A, 673, A148
- H. E. S. S. Collaboration et al. (2024) H. E. S. S. Collaboration, Aharonian, F., Ait Benkhali, F., et al. 2024, Science, 383, 402
- H. E. S. S. Collaboration et al. (2023b) H. E. S. S. Collaboration, Aharonian, F., Ait Benkhali, F., et al. 2023b, A&A, 672, A103
@@
- Lau et al. (2017) Lau, J. C., Rowell, G., Voisin, F., et al. 2017, PASA, 34, e064
- LHAASO collaboration (2021) LHAASO collaboration. 2021, arXiv e-prints, arXiv:2101.03508
@@
- Mohrmann et al. (2019) Mohrmann, L., Specovius, A., Tiziani, D., et al. 2019, A&A, 632, A72
@@
- Vovk et al. (2018) Vovk, I., Strzys, M., & Fruck, C. 2018, A&A, 619, A7
Appendix A Integral sensitivities
An often adopted approach in gamma-ray astronomy is to compare integral source fluxes and sensitivities, rather than differential ones. This technique is chosen when dealing with weak sources, falling off the stringent bin-per-bin criteria set in Sec. 4. The computation of integral sensitivity, in fact, requires that the very same criteria are satisfied above a given energy threshold, , namely that a certain number of events and an overall significance against background are obtained above . The integral method should, however, be used with caution because of three main reasons: i) a comparative analysis of instruments that perform their best in different energy ranges might be misleading; ii) its results depend on the spectral slope of the target source, particularly at low energy as shown later, implying that a rigorous comparison among source flux and sensitivity can only be performed when the two refer to the very same spectral slope; and iii) as in the integration one loses the information about the spectral features beyond , such as a cut-off, integral sensitivities might be overestimating the expected detection performance. In particular, this might be the case for LHAASO first catalogue sources, for which no information is currently available concerning the maximum photon energy detected for each of them. For these reasons, we adopted the differential approach as a benchmark.
Still, we find it worth to also discuss the results of the integral approach, as applied to extended sources, in order to understand its impact on source detectability. Being the criteria less stringent, a sensitivity improvement can in fact be expected, however the comparison with integrated fluxes yields a non trivial outcome, as we are going to discuss. Integral sensitivities for extended sources are provided in Fig. 11, as computed with the Crab-like spectrum defined in the text.
The impact of the chosen spectral slope, , both in the differential and in the integral sensitivity, , is shown in the left panel of Fig. 12 in terms of the ratio between calculated for CTA-South, and . Similar trends applies also to the other detectors.
As expected, differential sensitivities are not affected by the slope of the spectrum, while integral sensitivities are very sensitive to this choice, and the major discrepancies emerge below 1 TeV.
This result highlights that the differential approach is more preferable when discussing the detection prospects of sources with variegated spectral shapes, as it is the case of catalogues. On the other hand, if one focuses on the highest energies ( TeV), the dependence on the slope is almost negligible.
A proper comparison among differential and integral sensitivities can be achieved by multiplying the former to the central energy of the bin . The ratio of integral sensitivity, , to is shown in the right panel of Fig. 12, for different assumptions of source spectral slopes and for all considered detectors. The discussed slopes in this case coincide with the mean values of the distribution of slopes found for KM2A and HGPS sources, respectively and . We also consider the case for comparison. As it can be seen, the usage of integral sensitivities can actually provide an improvement compared to the differential sensitivities at VHEs by an amount that depends on both energy and spectral slope of the source. Around 50 TeV, which is the reference value for following-up KM2A sources, the improvement is assessed around 30-35% for IACTs. At lower energies, namely around 1 TeV, the improvement is slightly larger, up to factor 5, for the hardest considered spectra. However one should note that, as the integral sensitivity decreases, also the integral source flux gets reduced depending on its slope. In fact, the integral flux of a power-law function e.g. above a certain energy is
meaning that the source integral flux will be times smaller than its differential energy flux , for any . In particular, for the LHAASO-KM2A sources, this factor amounts to , while for HGPS sources it is . In both cases, as the flux reduction is more contained than the sensitivity gain, the integral approach actually provides a net gain in terms of detectability. Nonetheless, if compared at 50 TeV, the two effects are almost compensated for such that we do not envision an extension of the sample of KM2A sources that could be detected by IACTs, with respect to our results from Sec. 5.1 with the differential approach. Conversely, at 3 TeV, the larger sensitivity gain from the integral approach is not expected to enlarge the sample of detectable sources, as most of them were already shown to be in the reach of IACTs with the more stringent differential approach, rather to reduce the exposure time required.
Appendix B IACT worsening of IRFs with offset
@@
In the case of IACTs, extended source sensitivity is further degraded with respect to that presented in Sec. 4 if observations are carried out off-axis. Reduced performance are expected both in angular resolution and effective area. A preliminary estimation of the differential sensitivity degradation for extended source observations was performed in Ambrogi et al. (2018) for the previous configuration of CTA-South: a limited worsening was obtained for sources with radial extension below 2 deg (within a factor 2 for energies above 50 GeV). Nowadays, public IRFs are available including off-axis responses, such that more detailed evaluations can be performed. In Figs. 13, 14, and 15 we show respectively ASTRI, CTA-North, and CTA-South performances as a function of the offset angle of the observation with respect to the pointing direction of the telescope (fixed to zenith of 20 deg). For ASTRI, the impact of off-axis observations will mainly consist into a reduced effective area and enlarged angular resolution. For CTA, in addition, an increase in the energy threshold is expected.
@@
Appendix C IACT sensitivity for high zenith angle observations
IACT observations at high-zenith angles are characterised by a larger effective area at high energies at the expenses of an increased energy threshold. In fact, high-energy quasi-horizontal showers can profit of a better visibility, thanks to the larger portion of the atmosphere where they develop. On the other hand, low-energy high-zenith showers are less efficiently detected as their first interaction vertex occurs further away from the telescopes. As a result, the differential sensitivity of IACTs are affected by observations of sources at different zenith angles during the acquisition. We quantify the impact of this effect by computing the expected sensitivity for point-like observations of the CTA-N and CTA-S observatories, as shown in Fig. 17. As visible, an improvement of a factor 2 is expected above a few tens of TeV for CTA. We expect similar considerations to hold for ASTRI as well, however we cannot proceed to a proper evaluation of it since IRFs at high-zenith angle observations are not yet publicly available.
Appendix D Point-like differential sensitivities: Comparison with official results and investigation of CTA subarrays